# Logarithmic or algebraic: roughening of an active Kardar-Parisi-Zhang surface

### v3

@@ Abstract
We further find that the equal-time height-difference correlator $C_{h}(r,0)\equiv\langle[h({\bf x},t)-h(0,t)]^{2}\rangle\sim\frac{D_{0}}{\nu_{0}}[\ln(r/a)]^{\mu}$, for large $r\equiv|{\bf x}|\gg a$, indicating logarithmically faster or slower rise with the separation $r$ for large $r$ Haldar et al. (2023a, b), again generalising .the well-known QLRO found in an EW surface at 2D.
@@
Our continuously varying scaling exponents are a crucial outcome of the nonrenormalisation of $\lambda,\lambda_{1}$ and $\lambda_{2}$, rendering $\gamma_{1},\gamma_{2}$ marginal, which have been demonstrated at the one-loop order. Unlike the usual KPZ equation, Galilean invariance of the present model ensures nonrenormalisation of a combination of $\lambda,\,\lambda_{1},\,\lambda_{2}$, and not each of them individually. Thus, there is no surety that $\gamma_{1},\gamma_{2}$ should remain marginal even at higher-loop orders. We now argue that these possible higher-loop contributions, even though may exist, actually do not matter. Since we are at the critical dimension, $g(l)\sim 1/l$ for large $l$ at the one-loop order. At higher-loop orders, the Feynman diagrams will contain higher power of $g$. Hence, a general scaling solution for $g(l)$ should have the form $g(l)\sim 1/l+\sum_{n}c_{n}/l^{n},\,n>1$ is an integer. Thus, the higher-loop corrections to the one-loop solution of $g(l)$ should vanish like $1/l$ to a power greater than 1. Therefore, their integrals over $l$ from zero to infinity will be finite, so they will not change the anomalous behavior of $D$ and $\nu$. Similarly, they cannot make any divergent contribution to $\gamma_{1}(l)$ and $\gamma_{2}(l)$, even though there can be higher-loop diagrams. Therefore, our one-loop results are, in fact, asymptotically exact. This then implies that the continuous variation of the scaling exponents, making them nonuniversal, is also asymptotically exact in the long wavelength limit. See Refs. Haldar et al. (2023a, b); Sarkar and Basu (2012); Basu and Frey (2004, 2009); Bratanov et al. (2015); Mahault et al. (2018); Haldar and Basu (2022) for similar nonuniversal scaling exponents in other models.
@@
At higher dimensions $d>2$, the chiral term with coupling $\lambda_{2}$ cannot exist. The other two achiral nonlinear terms in Eq. (15) are present at $d>2$. The RG recursion relations for $d>2$ can be obtained from the Feynman diagrams given in SM with $\gamma_{2}=0$. Using a $d=2+\epsilon$ expansion as in the KPZ equation Tang et al. (1990), we find at the one-loop order or to the lowest order in $\epsilon$
@@
Parameter $\gamma_{1}$ remains marginal at the lowest order. Therefore, if ${\cal\tilde{A}}(\gamma_{1})>0$, $g(l)$ flows to zero rapidly, with $g(l)\sim g(0)\exp(-\epsilon\,l)$ in the long wavelength limit; $g^{*}=0$ is the only FP that is globally stable. This renders the nonlinearities irrelevant in the RG sense. Therefore, scaling in the long wavelength limit is identical to that in the EW equation: $z=2,\,\chi=(2-d)/2$. Furthermore, $d=2$ is then the upper critical dimension. On the other hand, if ${\cal\tilde{A}}(\gamma_{1})<0$, $g(l)$ has three FPs: $g^{*}_{c}=-\epsilon/{\cal\tilde{A}}(\gamma_{1})$, an unstable FP, parametrised by $\gamma_{1}$ and separating possibly two stable FPs, one being at $g^{*}=0$ Gaussian FP with EW scaling for that $d$, and another putative perturbatively inaccessible FP, corresponding presumably to an algebraically rough phase. This gives, with 2D as the lower critical dimension, a roughening transition at $d>2$, very similar to the KPZ equation at $d>2$, but with one caveat. At this unstable FP, using (3) and (4), to ${\cal O}(\epsilon)$ $z=2+\epsilon\frac{{\cal\tilde{C}}(\gamma_{1})}{{\cal\tilde{A}}(\gamma_{1})},\,\,\chi=-\epsilon\frac{{\cal\tilde{C}}(\gamma_{1})}{{\cal\tilde{A}}(\gamma_{1})}$, depend explicitly on $\gamma_{1}$ and deviate from their linear theory (or EW equation) values already at ${\cal O}(\epsilon)$. This is in contrast to the KPZ equation at $d>2$, where $z$ and $\chi$ at the unstable FP are at least ${\cal O}(\epsilon^{2})$ Tang et al. (1990). In fact, application of the Cole-Hopf transformation shows that $z=2,\,\chi=0$ at the unstable FP of the KPZ equation at $d>2$ Janssen et al. (1999).
@@
In the parameter space where ${\cal\tilde{A}}(\gamma_{1})<0$, solution of ${\cal\tilde{C}}(\gamma_{1})=0$ gives the red dashed lines $\gamma_{1}=\gamma^{-},\gamma^{{\dagger}}$ where, $\gamma^{{\dagger}}=0$ and $\gamma^{-}=-1.25$; see Fig. 2(a) for variation of $z$ and $\chi$ with $\gamma_{1}$ for a fixed $\epsilon$. Green strips correspond to ${\cal\tilde{C}}(\gamma_{1})>0$ where $\chi>0$ and $z<2$; ${\cal\tilde{C}}(\gamma_{1})<0$ is the blue region where, $\chi<0$ and $z>2$. For a given $\epsilon$, maximum value of $z$ and minimum value of $\chi$ are $z_{\text{max}}=2+0.292\epsilon$ and $\chi_{\text{min}}=-0.292\epsilon$ at $\gamma_{1}=-0.651$, such that the dynamics is slowest and the surface is smoothest at the unstable FP. Since $\chi_{\text{min}}>\chi_{\text{EW}}=-\epsilon/2$, an a-KPZ surface at the unstable FP is always rougher than an EW surface.
@@
Figure 2: (a) Variation of $z$ and $\chi$ with $\gamma_{1}$ on the fixed line $g_{c}^{*}=-1/{\cal\tilde{A}}(\gamma_{1})$ for $\epsilon=1$, (b) RG flow diagram in the $g-\gamma_{1}$ plane for $d>2$. The black dashed line is the fixed line $g_{c}^{*}=-1/{\cal\tilde{A}}(\gamma_{1})$ for $\epsilon=1$, bounded by lines $\gamma_{1}=\gamma_{-},\gamma_{+}$ (blue dashed lines). Stable (unstable) flow lines are the arrows pointing towards (away from) $g=0$ (see text).
@@
In the $g-\gamma_{1}$ plane, $g_{c}^{*}=-\epsilon/{\cal\tilde{A}}(\gamma_{1})$ is a fixed line, such that RG flow lines with initial $g$ values above the line flows to perturbatively inaccessible FP, see Fig. 2(b). And for systems with initial $g$ values lying below the line, the RG flow lines run parallel to the $g$-axis towards Gaussian FP, corresponding to the smooth phase belonging to the EW class. This behaviour holds within a range $\gamma_{-}>\gamma_{1}>\gamma_{+}$. As $\gamma_{1}\rightarrow\gamma_{+},\,\gamma_{-}$, ${\cal\tilde{A}}(\gamma_{1})$ vanishes and $g_{c}^{*}$ diverges. As soon as $\gamma_{1}$ exceeds $\gamma_{+}$ or falls short of $\gamma_{-}$, $g_{c}^{*}$ no longer exists with the roughening transition disappearing. RG flow lines starting from any initial condition with $\gamma_{1}>\gamma_{+}$ or $\gamma_{1}<\gamma_{-}$ (red region) where ${\cal\tilde{A}}(\gamma_{1})>0$, flow to $g^{*}=0$ ensuing scaling belonging to the EW class. Thus the roughening transition of the KPZ equation is entirely suppressed by the nonlocal nonlinear term.
@@
In summary, we have proposed and studied an “active KPZ” equation, having a surface velocity $v_{p}$ depending nonlocally on the surface gradients. Surprisingly, we find stable surfaces with positional generalised QLRO or generalised logarithmic roughness with nonuniversal exponents for wide-ranging choices of the model parameters, unlike the 2D KPZ equation. Physically, this is due to the competition between the nonlocal and local nonlinear terms and the lack of their renormalisation. At $d>2$, similar choices for the same parameters can suppress the KPZ roughening transitions, resulting into only smooth surfaces. Heuristically, a nonlocal part in $v_{p}$ means a local large fluctuation can generate a propulsion not just locally, but over large scales, which when sufficiently strong can suppress local variations in $v_{p}$ due to the local KPZ-nonlinear term. This in turn has the effect of reducing surface fluctuations. For other parameter choices, a KPZ-like perturbatively inaccessible rough phase is speculated. This may be explored by mode-coupling methods Bhattacharjee (1998). In that parameter space, the roughening transition survives at $d>2$, but with significantly different scaling properties, again with nonuniversal exponents. We hope our studies here will provide further impetus to study nonlocal effects on similar nonequilibrium surface dynamics models, e.g., the conserved KPZ Sun et al. (1989); Caballero et al. (2018) and the $|{\bf q}|$KPZ Besse et al. (2023) equations.
@@
Acknowledgement:- A.B. thanks the SERB, DST (India) for partial financial support through the MATRICS scheme [file no.: MTR/2020/000406].
Appendix A Active KPZ equation
@@
We first construct the appropriate equation in terms of the vector field ${\bf v}({\bf r},t)\equiv{\bm{\nabla}}h$. Field $\bf v$ has a definite physical interpretation: for an active membrane $\bf v$ is the fluctuation of the local normal to the membrane surface Chaikin and Lubensky (2000). Noting that the Burgers velocity $\bf v$ is a conserved vector field, it must follow a generic conservation law of the form
@@
where $J_{ij}$ is the velocity current. In general $J_{ij}$ can be decomposed as the sum of a symmetric part $J_{ij}^{s}=J_{ji}^{s}$ and an antisymmetric part $J_{ij}^{a}=-J_{ji}^{a}$: $J_{ij}=J_{ij}^{s}+J_{ij}^{a}$. Since $\bf v$ is purely irrolational, it can be expressed, via the Helmholtz theorem Bladel (1958); Koenigsberger (1906), solely in terms of its divergence, i.e., ${\bm{\nabla}}\cdot{\bf v}\equiv{\cal D}$, which follows the equation
@@
where ${\bm{\mu}}$ is the “chemical potential” vector, which in general can have irrolational and solenoidal parts. The latter part does not contribute to the dynamics of ${\cal D}$ and hence of $\bf v$, and hence can be set to zero, leaving $\bm{\mu}$ purely irrolational. In that case,
@@
Now consistent with the conservation law form of the Burgers equation, $\mu_{i}$ can be generalised from its form for the Burgers equation: It also admits, at the same order, a second contribution of the form $\sim\nabla_{j}(v_{i}v_{j})/\nabla^{2}$. Indeed, including this term the most general equation of $\bf v$ retaining only up to the lowest order in nonlinearities and spatial gradients that now includes a chiral contribution is of the form
@@
Both the additional nonlinear terms are nonlocal contributions to $\mu$. Equation (14) reduces to the well-known Burgers equation Forster et al. (1977) when $\lambda_{1}=0=\lambda_{2}$. Further, the tensor $e_{ij}$ is the 2D totally antisymmetric matrix, with $e_{11}=0=e_{22}$ and $e_{12}=1=-e_{21}$. Thus, the $\lambda_{2}$-term in Eq. (14) is the chiral contribution. Now, further demanding that ${\bf v}$ is fully irrotational with ${\bf v}={\bm{\nabla}}h$, the “superfluid velocity” of an XY model, or the deviation of the local normal of an interface in the Monge gauge Nelson et al. (1989), we obtain,
$+\lambda_{2}Q_{ij}({\bf r})e_{jm}(\nabla_{i}h\nabla_{m}h)+\eta,$
@@ Appendix B Galilean invariance
Using these we show below that if the height function $h({\bf x},t)$ transforms as $h^{\prime}({\bf x}^{\prime},t^{\prime})\rightarrow h({\bf x},t)+\textbf{c}\cdot\textbf{x}$, then Eq. (15) is invariant. Considering RHS of Eq. (15), the first term transforms as,
@@
$\lambda_{2}\frac{\nabla^{\prime}_{i}\nabla^{\prime}_{j}}{\nabla^{\prime 2}}e_{jm}(\nabla^{\prime}_{i}h^{\prime}\nabla^{\prime}_{m}h^{\prime})\rightarrow$ $\lambda_{2}\frac{\nabla_{i}\nabla_{j}}{\nabla^{2}}e_{jm}(\nabla_{i}h\nabla_{m}h)$
$-\lambda_{2}e_{ij}c_{i}\frac{\partial h}{\partial x_{j}}.$ (21)
Noise term remains invariant under this transformation. Thus after the transformation extra terms in RHS are
@@
Appendix C Renormalisation group calculations

### v4

@@ Abstract
The equal-time height-difference correlator $C_{h}(r,0)\equiv\langle[h({\bf x},t)-h({{\bf 0}},t)]^{2}\rangle\sim\frac{D_{0}}{\nu_{0}}[\ln(r/a)]^{\mu}$, for large $r\equiv|{\bf x}|\gg a$, indicating logarithmically faster or slower rise with the separation $r$ for large $r$ Haldar et al. (2023a, b), again generalizing the well-known QLRO found in a 2D EW surface.
@@
Our continuously varying scaling exponents are a crucial outcome of the nonrenormalization of $\lambda,\lambda_{1}$ and $\lambda_{2}$, rendering $\gamma_{1},\gamma_{2}$ marginal, which have been demonstrated at the one-loop order. Unlike the usual KPZ equation, Galilean invariance of the present model ensures nonrenormalization of a combination of $\lambda,\,\lambda_{1},\,\lambda_{2}$, and not each of them individually. Thus, there is no surety that $\gamma_{1},\gamma_{2}$ should remain marginal even at higher-loop orders. We now argue that these possible higher-loop contributions, even though they may exist, actually do not matter. For large $l$, $g(l)\sim 1/l$ at the one-loop order. At higher-loop orders, the Feynman diagrams will contain higher power of $g$. Hence, a general scaling solution for $g(l)$ should have the form $g(l)\sim 1/l+\sum_{n}c_{n}/l^{n},$ and $n>1$ is an integer. Thus, the higher-loop corrections to the one-loop solution of $g(l)$ should vanish like $1/l^{s},\,s>1$. Therefore, their integrals over $l$ from zero to infinity will be finite, so they will not change the anomalous behavior of $D$ and $\nu$. Similarly, they cannot make any divergent contribution to $\gamma_{1}(l)$ and $\gamma_{2}(l)$, even though there can be higher-loop diagrams. Therefore, our one-loop results are, in fact, asymptotically exact. This then implies that the continuous variation of the scaling exponents, making them nonuniversal, is also asymptotically exact in the long wavelength limit. See Refs. Haldar et al. (2023a, b); Sarkar and Basu (2012); Basu and Frey (2004, 2009); Bratanov et al. (2015); Mahault et al. (2018); Haldar and Basu (2022) for similar nonuniversal scaling exponents in other models.
@@
At higher dimensions $d>2$, the chiral term with coupling $\lambda_{2}$ cannot exist. The other two achiral nonlinear terms in Eq. (1) are present at $d>2$. The RG recursion relations for $d>2$ can be obtained from the Feynman diagrams given in the SM sup with $\gamma_{2}=0$. Using a $d=2+\epsilon$ expansion as in the KPZ equation Tang et al. (1990), we find at the one-loop order or to the lowest order in $\epsilon,$
@@
Parameter $\gamma_{1}$ remains marginal at the lowest order. Therefore, if ${\cal\tilde{A}}(\gamma_{1})>0$, $g(l)$ flows to zero rapidly, with $g(l)\sim g(0)\exp(-\epsilon\,l)$ in the long wavelength limit; $g^{*}=0$ is the only FP that is globally stable. This renders the nonlinearities irrelevant in the RG sense. Therefore, scaling in the long wavelength limit is identical to that in the EW equation: $z=2,\,\chi=(2-d)/2$. Furthermore, $d=2$ is then the upper critical dimension. On the other hand, if ${\cal\tilde{A}}(\gamma_{1})<0$, $g(l)$ has three FPs: $g^{*}_{c}=-\epsilon/{\cal\tilde{A}}(\gamma_{1})$, an unstable FP, parametrized by $\gamma_{1}$ and separating possibly two stable FPs, one being at $g^{*}=0$ Gaussian FP with EW scaling, and another putative perturbatively inaccessible FP, corresponding presumably to an algebraically rough phase. This gives, with 2D as the lower critical dimension, a roughening transition at $d>2$, very similar to the KPZ equation at $d>2$, but with one caveat. At this unstable FP, using (3) and (4), to ${\cal O}(\epsilon)\,$ ${z}=2+\epsilon\frac{{\cal\tilde{C}}(\gamma_{1})}{{\cal\tilde{A}}(\gamma_{1})},\,\,\chi=-\epsilon\frac{{\cal\tilde{C}}(\gamma_{1})}{{\cal\tilde{A}}(\gamma_{1})}$, depend explicitly on $\gamma_{1}$ and deviate from their linear theory (or EW equation) values already at ${\cal O}(\epsilon)$. This is in contrast to the KPZ equation at $d>2$, where $z$ and $\chi$ at the unstable FP are at least ${\cal O}(\epsilon^{2})$ Tang et al. (1990). In fact, application of the Cole-Hopf transformation shows that $z=2,\,\chi=0$ at the unstable FP of the KPZ equation at $d>2$ Janssen et al. (1999).
@@
When ${\cal\tilde{A}}(\gamma_{1})<0$, the solution of ${\cal\tilde{C}}(\gamma_{1})=0$ gives the red dashed lines $\gamma_{1}=\gamma^{-},\gamma^{{\dagger}}$ where, $\gamma^{{\dagger}}=0$ and $\gamma^{-}=-1.25$; see Fig. 2(a) for a variation of $z$ and $\chi$ with $\gamma_{1}$ for a fixed $\epsilon$. Green strips correspond to ${\cal\tilde{C}}(\gamma_{1})>0$ where $\chi>0$ and $z<2$; ${\cal\tilde{C}}(\gamma_{1})<0$ is the blue region where $\chi<0$ and $z>2$. For a given $\epsilon$, maximum value of $z$ and minimum value of $\chi$ are $z_{\text{max}}=2+0.292\epsilon$ and $\chi_{\text{min}}=-0.292\epsilon$ at $\gamma_{1}=-0.651$, such that the dynamics is slowest and the surface is smoothest at the unstable FP. Since $\chi_{\text{min}}>\chi_{\text{EW}}=-\epsilon/2$, an a-KPZ surface at the unstable FP is always rougher than an EW surface.
@@
Figure 2: (a) Variation of $z$ and $\chi$ with $\gamma_{1}$ on the fixed line $g_{c}^{*}=-1/{\cal\tilde{A}}(\gamma_{1})$ for $\epsilon=1.$ (b) RG flow diagram in the $g$-$\gamma_{1}$ plane for $d>2$. The black dashed line is the fixed line $g_{c}^{*}=-1/{\cal\tilde{A}}(\gamma_{1})$ for $\epsilon=1$, bounded by lines $\gamma_{1}=\gamma_{-},\gamma_{+}$ (blue dashed lines). Stable (unstable) flow lines are the arrows pointing toward (away from) $g=0$ (see text).
@@
In the $g$-$\gamma_{1}$ plane, $g_{c}^{*}=-\epsilon/{\cal\tilde{A}}(\gamma_{1})$ is a fixed line, such that RG flow lines with initial $g$ values above the line flows to perturbatively inaccessible FP, see Fig. 2(b). And for systems with initial $g$ values lying below the line, the RG flow lines run parallel to the $g$ axis toward Gaussian FP, corresponding to the smooth phase belonging to the EW class. This behavior holds within a range $\gamma_{-}>\gamma_{1}>\gamma_{+}$. As $\gamma_{1}\rightarrow\gamma_{+},\,\gamma_{-}$, ${\cal\tilde{A}}(\gamma_{1})$ vanishes and $g_{c}^{*}$ diverges. As soon as $\gamma_{1}$ exceeds $\gamma_{+}$ or falls short of $\gamma_{-}$, $g_{c}^{*}$ no longer exists with the roughening transition disappearing. RG flow lines starting from any initial condition with $\gamma_{1}>\gamma_{+}$ or $\gamma_{1}<\gamma_{-}$ (red region) where ${\cal\tilde{A}}(\gamma_{1})>0$, flow to $g^{*}=0$ ensuing scaling belonging to the EW class.
@@
In summary, we have proposed and studied an “active KPZ” equation, having a surface velocity $v_{p}$ depending nonlocally on the surface gradients. Surprisingly, we find stable surfaces with positional generalized QLRO or generalized logarithmic roughness with nonuniversal exponents for wide-ranging choices of the model parameters, unlike the 2D KPZ equation. Physically, this is due to the competition between the nonlocal and local nonlinear terms and the lack of their renormalization. Indeed, this competition between the nonlocal and local nonlinear terms distinguishes our model (1) from that studied in Ref. Mukherji and Bhattacharjee (1997), giving either generalized QLRO with nonuniversal scaling exponents or a novel roughening transition even in 2D controlled by the relative strengths of the local and nonlocal interactions. At $d>2$, sufficiently strong nonlinear nonlocal effects can either entirely suppress the KPZ roughening transitions, resulting into only smooth surfaces, or else give a roughening transition with nonuniversal scaling very different from the well-known roughening transition in the KPZ equation. Heuristically, a nonlocal part in $v_{p}$ means a local large fluctuation can generate a propulsion not just locally, but over large scales, which when sufficiently strong can suppress local variations in $v_{p}$ due to the local KPZ-nonlinear term. This in turn has the effect of reducing surface fluctuations. For other parameter choices, a KPZ-like perturbatively inaccessible rough phase is speculated. This may be explored by mode-coupling methods Bhattacharjee (1998). In that parameter space, the roughening transition survives at $d>2$, but with significantly different scaling properties, again with nonuniversal exponents. We hope our studies here will provide further impetus to study nonlocal effects on similar nonequilibrium surface dynamics models, e.g., the conserved KPZ Sun et al. (1989); Caballero et al. (2018) and the $|{\bf q}|$KPZ Besse et al. (2023) equations.
@@
A.B. thanks the SERB, DST (India) for partial financial support through the MATRICS scheme [File No. MTR/2020/000406].
Supplemental Material
I Active KPZ equation
@@
We first construct the active KPZ equation in terms of the vector field ${\bf v}({\bf r},t)\equiv{\bm{\nabla}}h$. Field $\bf v$ has a definite physical interpretation: for an active membrane $\bf v$ is the (small) fluctuation of the local normal to the membrane surface measured in the Monge gauge Nelson et al. (1989); Chaikin and Lubensky (2000). Noting that the Burgers velocity $\bf v$ is a conserved vector field, it must follow a generic conservation law of the form
@@
where $J_{ij}$ is the velocity current. In general $J_{ij}$ can be decomposed as the sum of a symmetric part $J_{ij}^{s}=J_{ji}^{s}$ and an antisymmetric part $J_{ij}^{a}=-J_{ji}^{a}$: $J_{ij}=J_{ij}^{s}+J_{ij}^{a}$. Since $\bf v$ is purely irrotational, it can be expressed, via the Helmholtz theorem Bladel (1958); Koenigsberger (1906), solely in terms of its divergence, i.e., ${\bm{\nabla}}\cdot{\bf v}\equiv{\cal D}$, which follows the equation
@@
where $\bm{\mu}$} is the “chemical potential” vector, which in general can have irrotational and solenoidal parts. The latter part does not contribute to the dynamics of $\cal D$} and hence of $\bf v$, and hence can be set to zero, leaving $\bm{\mu}$ purely irrotational. In that case,
@@
Now consistent with the conservation law form of the Burgers equation, $\mu_{i}$ in the Burgers equation can be generalised. It also admits, at the same order, a second contribution of the form $\sim\nabla_{j}(v_{i}v_{j})/\nabla^{2}$. Indeed, including this term the most general equation of $\bf v$ retaining only up to the lowest order in nonlinearities and spatial gradients that now includes a chiral contribution is of the form
@@
Both the additional nonlinear terms make nonlocal contributions to $\mu$. Equation (14) reduces to the well-known Burgers equation Forster et al. (1977) when $\lambda_{1}=0=\lambda_{2}$. Further, the tensor $e_{ij}$ is the 2D totally antisymmetric matrix, with $e_{11}=0=e_{22}$ and $e_{12}=1=-e_{21}$. Thus, the $\lambda_{2}$-term in Eq. (14) is the chiral contribution. Now, further demanding that $\bf v$} is fully irrotational with ${\bf v}={\bm{\nabla}}h$, the “superfluid velocity” of an XY model, or the deviation of the local normal of an interface in the Monge gauge Nelson et al. (1989), we obtain,
$\frac{\partial h}{\partial t}=$ $\nu\nabla^{2}h+\frac{\lambda}{2}(\bm{\nabla}h)^{2}+\lambda_{1}Q_{ij}({\bf r})(\nabla_{i}h\nabla_{j}h)+\lambda_{2}Q_{ij}({\bf r})e_{jm}(\nabla_{i}h\nabla_{m}h)+\eta,$ (15)
@@ II Galilean invariance
Using these we show below that if the height function $h({\bf x},t)$ transforms as $h^{\prime}({\bf x}^{\prime},t^{\prime})\rightarrow h({\bf x},t)+\textbf{c}\cdot\textbf{x}$, then Eq. (1) remains invariant. Considering RHS of Eq. (1), the first term transforms as,
@@
$\lambda_{2}\frac{\nabla^{\prime}_{i}\nabla^{\prime}_{j}}{\nabla^{\prime 2}}e_{jm}(\nabla^{\prime}_{i}h^{\prime}\nabla^{\prime}_{m}h^{\prime})\rightarrow\lambda_{2}\frac{\nabla_{i}\nabla_{j}}{\nabla^{2}}e_{jm}(\nabla_{i}h\nabla_{m}h)-\lambda_{2}e_{ij}c_{i}\frac{\partial h}{\partial x_{j}}.$ (21)
Noise term remains invariant under this transformation. After the transformation, the extra terms in RHS are
@@
III Renormalisation group calculations